Please, this is urgent , solve it as soon as possible.

Suppose V is finite-dimensional and phi is a linear functional on V. Then
there is a unique vector u in V such that
phi(v)=< v, u> for every v in V.
1
Expert's answer
2020-06-15T18:57:28-0400

The vector "u"  is unique. Let "u_1"  and "u_2"  be such that "\phi(v) = \langle v, u_1\rangle"  and "\phi(v) = \langle v, u_2\rangle"  for every "v\in V". Then "\langle v, u_1-u_2\rangle = \langle v, u_1\rangle - \langle v, u_2\rangle = \phi(v) - \phi(v) = 0"  for every "v\in V" . Substituting "u_1-u_2"  for "v" , we get "\langle u_1-u_2, u_1-u_2\rangle = 0" , hence "u_1-u_2 = 0" .




The vector "u"  exists. Since "V" is finite-dimensional, the Gram-Schmidt procedure gives an orthonormal basis "\{e_i\}_i"  of "V". Choose "u = \sum_i \overline{\phi(e_i)} e_i" .


Let "v\in V". Since "\{e_i\}_i"  is a basis, there are scalars "\{a_i\}_i"  such that "v = \sum_i a_i e_i" . Hence


"\phi(v) = \phi\left(\sum_i a_i e_i\right) = \sum_i a_i \phi(e_i)."

On the other hand,


"\langle v, u\rangle = \left\langle \sum_i a_i e_i, \sum_i \overline{\phi(e_i)} e_i\right\rangle \\= \sum_i \sum_j a_i \phi(e_j) \langle e_i, e_j\rangle = \sum_i a_i \phi(e_i)"

where the last equality follows from the orthonormality of "\{e_i\}_i" . Therefore, "\phi(v) = \langle v, u\rangle" for every "v\in V".


The expression "\overline{a}"  denotes the complex conjugate of a complex scalar "a"  and is used for complex vector spaces. For real vector spaces, the proof is the same as above except that complex conjugation should be eliminated.